OpenStudy (anonymous):

If a snowball melts so that its surface area decreases at a rate of 9 cm2/min, find the rate at which the diameter decreases when the diameter is 12 cm.

zepdrix (zepdrix):

Hmmmmmmmmmmmmmm So we'll assume the snowball is spherical :) The Surface Area of a sphere is given as: \[\huge A=4 \pi r^2\] I think.. right?

zepdrix (zepdrix):

They gave us some information.. they told us the instantaneous rate of change of the Surface Area: \[\huge A'=-9cm^2/\min\]

zepdrix (zepdrix):

Negative since it's DECREASING! :O

zepdrix (zepdrix):

So they told us that when the diameter (d) is a particular value, find d'

zepdrix (zepdrix):

So we want to rewrite our Area formula is terms of DIAMETER, not RADIUS. So how can we replace the r^2 with d's?

OpenStudy (anonymous):

totally missed that part^

OpenStudy (anonymous):

r = d/2

OpenStudy (anonymous):

since you have A' you can use that to get d' im guessing

zepdrix (zepdrix):

\[\huge A=4\pi(r)^2=4\pi\left(\frac{ d }{ 2 }\right)^2\]\[\huge =\pi(d)^2\]

zepdrix (zepdrix):

Yes! true story! :D

OpenStudy (anonymous):

a' = 2pi d

zepdrix (zepdrix):

Just don't forget to square the entire substitution you make when you plug in the d/2! the 1/2 get's squared also! :D cancelling out with the 4.

OpenStudy (anonymous):

-9/2pi = d

zepdrix (zepdrix):

Remember, we're differentiating with respect to TIME. so not only do we get an A', but we should also get a d' when we differentiate d.

OpenStudy (anonymous):

uhm. yes..

OpenStudy (anonymous):

so do you just get 2d * d' in this case

OpenStudy (anonymous):

and then divide it.. because it's given?

zepdrix (zepdrix):

yes looks good! c: And from that point, we're left with 3 variables (2 of which they TOLD US!) and we can solve for d' from there! :D

OpenStudy (anonymous):

it's.. magic.

zepdrix (zepdrix):

hah XD you know it cool cat!

OpenStudy (anonymous):

you should get a trophy

zepdrix (zepdrix):

Pshhhh i got a medal :D good enough!